Circular No. 6311

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) D. Schleicher, Lowell Observatory, reports: “Narrow-band photometry

of this comet was obtained using the Lowell Observatory 0.8-m telescope on Feb. 9.5 UT. Haser-model production rates were: log Q(OH) = 28.70, log Q(CN) = 26.16, and log Q(C_{2}) = 26.15. The dust production rate was log (Afρ) = 3.19. For comparison, the water production rate (based on OH) is ∼ 30 percent less than we observed for 1P/Halley at a comparable heliocentric distance.”

Total-magnitude and coma-diameter estimates (cf. IAUC 6304): Jan. 1.84 UT, 13.3, 1′ (K. Takamizawa, Saku, Nagano, Japan, 0.10-m refractor; prediscovery photographs; low altitude); Feb. 10.76, 8.7, 3′ (T. Lovejoy, Jimboomba, Queensland, 0.25-m reflector; visual).

If the Jan. 1 positions can be relied on to arcsecond accuracy, the ephemeris uncertainty in late March is at most a few tenths of a degree. Improved ephemeris from orbital elements by B. G. Marsden on MPEC 1996-C06 (minimum distance from the earth is 0.10 AU on Mar. 25):

1996 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Feb. 7 14^{h}36.^{m}95 −24^{o}54.^{′}8 1.592 1.914 92.^{o}9 31.^{o}0 9.3
12 14 40.64 −24 47.6 1.424 1.831 97.1 32.3 8.9
17 14 44.10 −24 32.4 1.256 1.746 101.4  33.7 8.4
22 14 47.28 −24 05.8 1.088 1.659 105.9  35.0 7.9
27 14 50.11 −23 22.1 0.920 1.570 110.5  36.2 7.3
Mar. 3 14 52.52 −22 10.5 0.753 1.479 115.4  37.3 6.6
8 14 54.36 −20 08.7 0.586 1.386 120.6  38.0 5.8
13 14 55.36 −16 20.7 0.422 1.291 126.6  38.2 4.7
18 14 54.70 − 7 39.0 0.263 1.192 133.6  37.2 3.4
23 14 48.03 +22 34.8 0.127 1.090 134.8  40.4 1.4
28 4 05.6 +78 49.1 0.135 0.985 80.5 91.7 1.1
Apr. 2 3 11.39 +52 00.4 0.273 0.876  55.8 109.3 2.1
7 3 05.25 +43 28.1 0.429 0.762  45.1 111.3 2.5
12 3 00.61 +39 02.9 0.590 0.643  37.4 108.7 2.4
17 2 55.04 +35 47.4 0.754 0.519  30.3 102.6 2.0
22 2 47.46 +32 29.3 0.919 0.393  23.0  90.9 1.3
27 2 37.08 +27 59.6 1.081 0.278  14.7  67.2 0.1
1996 February 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green